The rate of the chemical reaction doubles for an increase of 10K in absolute temperature from 298K. Calculate Ea?

Grade:8

1 Answers

Sunil Kumar FP
9 years ago
The rate of the chemical reaction doubles for an increase of 10K in absolute temperature from 298K
log(k2/k1)=Ea/2.303 R(1/T1-1/T2)
it inpleis from the above value that
log2=Ea/2.303*8.314(1/298-1/308)
Ea=53KJ
The energy of activation is the minimum energy for a reaction to have in order to proceed in a given direction